OpenStudy (anonymous):

Identify the center and the radius of the circle: (x-3)^2+(y+1)^2=36 Thanks!

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf (x-h)^2+(y-k)^2=r^2 }\) \(\Large\color{blue}{ \bf center~~is~~(h,k)}\)

OpenStudy (solomonzelman):

hint \(\Large\color{red}{ \bf 36=6^2~~~and~~~1=-(-1) }\)

OpenStudy (solomonzelman):

what do you think about \(\Large\color{red}{ \bf (x-\color{blue} { 3 })^2+(y-\color{blue} { (-1)})^2=\color{blue} { 6 }^2 }\)

OpenStudy (anonymous):

um. The center is (3,-1)?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

ok cool! How do I find the radius?